Circular No. 7516

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE B. Schmidt, Australian National University, on behalf of the High-

Z Supernova Search Team (cf. IAUC 6646), reports the discovery of 11 supernovae on R-band CCD images taken with the Canada-France-Hawaii Telescope by J. Tonry, B. Barris, P. Garnavich, S. Holland, A. Riess, and himself; and on CCD images taken with the Cerro Tololo Interamerican Observatory Blanco telescope by N. B. Suntzeff, R. A. Schommer, R. C. Smith, P. Candia, P. Challis, S. Jha, L. Strolger, and A. Clocchiatti. The supernovae were identified by subtracting images taken in Sept. from those obtained in Oct., and confirmed with spectra and images taken on Oct. 31 and Nov. 1 with the European Southern Observatory 8.2-m ANTU Very Large Telescope (+ FORS1) by B. Leibundgut and J. Sollerman.

SN  2000 UT  α_{2000}  δ_{2000}  R  z Type
2000dy Oct. 25 23^{h}25^{m}35.^{s}93 − 0^{o}22′34′′.0 22.7 0.61 Ia
2000dz Oct. 21 23 30 41.36 + 0 18 42.7 23.1 0.50 Ia
2000ea Oct. 25 2 09 54.02 − 5 28 17.8 23.3 0.42 Ia
2000eb Oct. 25 2 10 25.75 − 5 25 33.7 23.6 0.68 Ia
2000ec Oct. 21 2 11 32.03 − 4 13 56.1 22.7 0.47 Ia
2000ed Oct. 21 2 11 06.91 − 4 22 38.6 23.7 0.41 II?
2000ee Oct. 21 2 27 34.53 + 1 11 49.4 22.6 0.47 Ia
2000ef Oct. 21 2 28 10.30 − 1 11 03.6 23.2 0.41 II?
2000eg Oct. 21 2 30 21.05 + 1 03 48.5 22.5 0.54 Ia
2000eh Oct. 21 4 15 02.44 + 4 23 18.1 22.4 0.49 Ia
2000ei Oct. 21 4 17 07.18 + 5 45 53.1 22.8 0.60 II?

The supernovae are all within 2′′ of their host galaxies. Objects marked with “II?” are dominated by a blue continuum and are not associated with their host galaxy nucleus. Finder charts are available for these objects

at http://oir-www.harvard.edu/cfa/oir/Research/supernova, as are

also information and finder charts for an additional 50 supernova candidates in the range R = 20–25 that do not have spectroscopic confirmation.

NOVA IN M31 L. Donato, S. Garzia, V. Gonano, and G. Sostero, Remanzacco Ob-

servatory, report their discovery of an apparent nova at α = 0^{h}42^{m}57^{s}.17, δ = +41^{o}07′15′′.4 (equinox 2000.0; measurement by K. Korlević at Višnjan). Unfiltered CCD magnitudes by Donato et al.: Sept. 22 UT, [18.5; 22.82, 17.6; 28.85, 18.1; Nov. 1.84, 18.7.

2000 November 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green